OpenStudy (paperbacon):

Need someones help with Algebra 2 asap! TIA! :)

OpenStudy (paperbacon):

Will fan and medal anyone that helps!

OpenStudy (paperbacon):

Can you take the square root of each side of an equation? This being said (x+5)^2+(y-7)^2 and (x+5)+(y-7) are equal?

OpenStudy (paperbacon):

@TheSmartOne @Miracrown

OpenStudy (paperbacon):

@Bee_see

OpenStudy (amistre64):

no: sqrt(a^2+b^2) does not necessarily equate to a+b

OpenStudy (amistre64):

proof let sqrt(a^2+b^2) = a+b square both side a^2 + b^2 = a^2 + b^2 +2ab 0 = 2ab only of a or b or both equal 0